OpenStudy (moirrraaa):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (moirrraaa):

f(x)= -2(x-4)^2=2 g(x)= 5x^2-10x+7

OpenStudy (welshfella):

f(x) is witten in vertex form a(x - b)^2 + c where axis of symmetry is x = b

OpenStudy (welshfella):

to find the axis of symmetry of g(x) convert it to vertex form and use the above rule

OpenStudy (welshfella):

foor h(x) the axis of symmetry is the equation of the line passing therough the vertex

OpenStudy (welshfella):

has this helped?